HUBBLE SPACE TELESCOPE DAILY REPORT #1863 PERIOD COVERED: 0000Z (UTC) 04/23/97 - 0000Z (UTC) 04/24/97 Daily Status Report as of 114/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed NIC/3 7043 (NICMOS Focus Monitor) The Near Infrared Camera (Camera 3) was used to make observations = of NGC-3603 in order to make focus calibration and monitoring = measurements. This proposal is built to run NIC/3 focus sweeps twice a = week during SMOV. The observations were executed as scheduled, and no = anomalies were noted. 1.2 Completed Four Sets of WF/PC-2 7553 (Pure Parallel Archive Survey = II) WF/PC-2 parallel observations were obtained in the ultraviolet in = search for interesting objects. This program is designed to find = extragalactic sources with near ultraviolet excesses (3000 A0) outside = the galactic plane and far ultraviolet sources (1600 A0) in the galactic = plane. The procedure was executed as scheduled, and no anomalies were = reported. 1.3 Completed NIC /3 7055 (NICMOS Grism Validation) The Near Infrared Camera #3 was used to make observations of the = object K4-47 in order to perform validation of the operations of the = three NICMOS grisms and supply the initial level of calibration for = spectroscopic observations. The observations were completed as planned = and no problems were reported. 1.4 Completed FOC/96 6894 (Cycle 6: Aperture Monitor Location) The Faint Object Camera (f/96) was used to make observations of = selected targets in NGC-188 in order to monitor the location of the = camera's aperture with respect to the V2-V3 coordinate frame. The = observations were completed as planned, and no problems were reported. 1.5 Completed WF/PC-2 6480 (The Butcher-Oemler Effect: Where Have All = The Spirals Gone?) The WF/PC-2 was used to make observations of the blue disk-type = galaxy A851-POS1 in order to study the Butcher-Oemler effect. Other = observational data show that mergers and tidal interactions were apparentl= y more common in younger clusters; these may contribute to the = disappearance of spiral galaxies by the present epoch. The observations = were executed as scheduled, and no anomalies were noted. 1.6 Completed WF/PC-2 6902 (WF/PC-2 Cycle 6 Photometric Calibration = Monitor UV/OPT STD) The WF/PC-2 was used to make both ultraviolet and optical = photometric observations of the standard star GRW+70D582 as part of a = Cycle 6 calibration procedure. This observations was done immediately = before and after decontamination procedures. The observations were = completed as planned, and no problems were reported. 1.7 Completed NIC/1/2/3 7051 (NICMOS Detector Noise and Dark = Characterization - Non-Occulted) The Near Infrared Cameras (NIC 1, 2 and 3) were used to detector = noise and dark characterizational calibration observations. The purpose = of this proposal is to characterize NICMOS read noise and dark current = for the three cameras, verification of the stability since SLTV and = measurements of the cosmic ray event outside SAA. The observations were = completed as planned, and no problems were reported. 1.8 Completed NIC/1/2/3 7048 (NICMOS Parallel Thermal Background Test) The Near Infrared Cameras (Cameras 1, 2 and 3) were used to make = observations of an object designated as Target-1 in order execute = measurements of the stability of the HST thermal emissions. The = observations were executed as scheduled, and no problems were reported. 1.9 Completed NIC/2 7154 (NICMOS Coronographic Hole Monitor) The Near Infrared Camera (NIC 2) was used to make observations of = object NGC-3606 in order to monitor the effects of coronagrapghic hole. = The purpose of this proposal is to take lamp flats with camera 2 to = monitor the location of the coronographic spot. The observations were = completed as scheduled, and no anomalies were reported. 1.10 Completed NIC/1/2/3 7045 (NICMOS Persistence Test) The Near Infrared Cameras (NIC 1, 2 and 3) were used to make = observations of standard star HD106965 in order to measure the system's = persistence. This test measures the effects of severe overexposure of = the NICMOS detectors to determine if operational limitations are required = for programs expecting such over exposures. It also obtained data on the = extreme wings of the NICMOS PSF. Dark calibration observations were also = taken. The guide star acquisition for this proposal (at 113/232959Z) = defaulted to fine lock on FGS #2 when the search limit was exceeded on = FGS #3. HSTAR 6044 was written. The observations were completed as = planned, and no affects on science were reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 4 Successful : 4 As described in 1.10, one acquisition defaulted to fine lock on = one FGS only. Scheduled Reacquisitions : 4 Successful : 4 2.2 FHST Updates: Scheduled : 8 Successful : 8 2.3 Operations Notes: The battery #2 capacity check is continuing. Four gyro configuation was successfully commanded at 113/2147Z and = the associated PRS operational limits were adjusted. STIS flight software version 4.6 was installed, with the MAMA 1 = and 2 off, at 113/2313Z. STIS was suspended at 113/2314Z. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science and SMOV operations. /DMH